5.Let A be an arbitrary vector and let n be a unit vector in some fixed direction. Show that A=(A*widehat n)widehat n+(widehat n times A)times widehat n ..
1
Expert's answer
2020-12-10T11:06:49-0500

Explanation & proof


  • The vector equation given in the description is

"\qquad\qquad
\small A=(A*\widehat n)\widehat n+(\widehat n \times A)
\times \widehat n"

  • The first term is some product of the three vectors where as the second is a vector product of the three.
  • In general, if a, b & c are three non-coplanar vectors then,

"\qquad\qquad
\begin{aligned}
\small \widehat a\times (\widehat b\times \widehat c)&= \small (\widehat a.\widehat c)\widehat b-(\widehat a.\widehat b)\widehat c
\end{aligned}" & "\qquad\qquad
\begin{aligned}
\small \widehat a\times (\widehat b\times \widehat c) = -(\widehat b\times \widehat c)\times \widehat a
\end{aligned}"

and "\small \widehat a.\widehat b=\widehat b.\widehat a"

  • Therefore, by applying these to the given one

"\qquad\qquad
\begin{aligned}
\small \widehat A &= \small (\widehat A*\widehat n)\widehat n+(\widehat n \times \widehat A)\times \widehat n\\
&= \small (\widehat A.\widehat n)\widehat n-\widehat n\times (\widehat n \times \widehat A)\\
&= \small (\widehat A.\widehat n)\widehat n-\big[(\widehat n.\widehat A)\widehat n-(\widehat n.\widehat n)\widehat A \big]\\
&= \small \cancel{(\widehat A.\widehat n)\widehat n}-\cancel{(\widehat A.\widehat n)\widehat n}+(\widehat n.\widehat n)\widehat A\\
&= \small (1)\widehat A\\
&= \small \widehat A
\end{aligned}"